# Regimes and Transitions of the Nonlinear Temporal Talbot Effect: Underlying Mechanism for A-Type Breathers, Soliton Crystals, and Soliton Gas

### v1

@@ 5 Conclusion
In Ref.  [60], my colleague Michael Böhm and I reported a strong relationship between the temporal Talbot effect and MI-related nonlinear-wave generation from a phase-modulated CW laser field in optical fibers. The results were achieved using the numerical technique of the Soliton Radiation Beat Analysis [61]. Thus, we reported input-power-dependent transitions from a quasi-linear regime to the nonlinear regimes of the nonlinear Talbot effect that include soliton crystals and separated solitons (also called soliton gas due to a weak interaction between each other). The transitions occurred at well-defined input power values. To the best of my knowledge, Ref. [60] reports the first results that discuss the nonlinear temporal Talbot effect in such detail. However, more studies were needed to develop an exhaustive theory of this effect.
Here, I build upon the results of Ref. [60] and study the transitions of the temporal Talbot effect and their physical mechanisms using Soliton Radiation Beat Analysis and deriving a dispersion relation for frequency-comb lines. The following results were achieved:
  1. i)
@@
  5. v)
    I show that the temporal Talbot effect determines the spatio-temporal distribution of the optical field by driving the (higher-order) spatial-mode interference and, thus, constitutes an underlying effect for optical-field evolution in the linear and nonlinear regimes. In particular, the spatial-mode interference relates to the the Fermi-Pasta-Ulam-Tsingou recurrence of optical pulses, an observation that needs to be studies in more detail for an exhaustive theory of the nonlinear temporal Talbot effect.
@@
To illustrate the applicability of this study to real-world applications, let us consider frequency-multiplexed optical computing. This study provides an answer to the question of what are the primary driving mechanisms of data processing (FWM or solitons?) in a frequency-multiplexed Extreme Learning Machine and explain the degradation of its performance at higher input powers as reported in Refs. [5], [6]. When a frequency comb is used as an information carrier, data processing occurs through frequency-comb lines influencing each other, which happens via FWM [6] or, as shown here, XPM. As discussed in this study, the regime of separated solitons primarily relates to SPM, i.e. self-influencing, of frequency comb lines, while the information processing effects of FWM and XPM are negligible. As a result, the performance of a frequency-multiplexed Extreme Learning Machine deteriorates in the regime of separated solitons which happens at higher input powers. Thus, FWM and XPM, rather than solitons, should be considered primary mechanisms in frequency-multiplexed optical computing. With it, an important question related to the system design and performance optimization of frequency-multiplexed optical computing schemes is solved.
@@ Acknowledgments
I would like to say thank you to Michael Böhm (Wismar University of Applied Sciences), the inventor of Soliton Radiation Beat Analysis, for our long and inspiring conversations about optical solitons.
Disclosures
I declare no conflict of interest.
Data availability

### v2

@@ 5 Conclusion
In Ref.  [60], my colleague Michael Böhm and I reported a strong relationship between the temporal Talbot effect and MI-related nonlinear-wave generation from a phase-modulated CW laser field in optical fibers. The results were achieved using the numerical technique of the Soliton Radiation Beat Analysis [61]. Thus, we reported input-power-dependent transitions from a quasi-linear regime to the nonlinear regimes of the nonlinear Talbot effect that include soliton crystals and separated solitons (also called soliton gas due to a weak interaction between individual pulses). The transitions occurred at well-defined input power values. To the best of my knowledge, Ref. [60] reports the first results that discuss the nonlinear temporal Talbot effect in such detail. However, more studies were needed to better explain the observations presented in Ref. [60] and to develop a more profound theory of the nonlinear temporal Talbot effect.
Here, I build upon the results of Ref. [60] and study the transitions of the temporal Talbot effect and their physical mechanisms using Soliton Radiation Beat Analysis and deriving a dispersion relation for frequency-comb lines. This approach allows me to incorporate nonlinear effects in the framework of the initially linear temporal Talbot effect, paving the way for the development of a concise theory of the nonlinear temporal Talbot effect. It offers a new perspective and understanding of nonlinear wave generation in dispersion media. Here, I concentrate on optical fibers, but the results are also transferable to semiconductor waveguides (cf. [26], [27]).
The following was achieved:
  1. 1.
@@
  5. 5.
    Most importantly, I show that the temporal Talbot effect determines the spatio-temporal distribution of the optical field by driving the (higher-order) spatial-mode interference and, thus, constitutes an underlying effect for optical-field evolution in the linear and nonlinear regimes. In particular, the spatial-mode interference relates to the the Fermi-Pasta-Ulam-Tsingou recurrence of optical pulses, an observation that needs to be studies in more detail for an exhaustive theory of the nonlinear temporal Talbot effect.
@@
The results of this study advance the fields of Nonlinear Optics and Wave Theory and are directly applicable to frequency-comb engineering and wave shaping, high-resolution spectroscopy, and optical systems, fibers (cf. [20], [21], [55], [57], [60]) and semiconductors (cf. Refs. [26], [27]), where controlling of nonlinear dynamics is critical.
To illustrate the applicability of this study to an application, let us consider frequency-multiplexed optical computing. This study provides an answer to the question of what the primary driving mechanisms of data processing (FWM or solitons?) are in a frequency-multiplexed Extreme Learning Machine and explains the degradation of its performance at higher input powers as reported in Refs. [5], [6]. When a frequency comb is used as an information carrier, data processing occurs through frequency-comb lines influencing each other, which happens via FWM [6] or, as shown here, XPM. As discussed in this study, the regime of separated solitons primarily relates to SPM, i.e., self-influencing, of frequency comb lines, while the information processing effects of FWM and XPM are negligible. As a result, the performance of a frequency-multiplexed Extreme Learning Machine deteriorates in the regime of separated solitons, which happens at higher input powers. Thus, FWM and XPM, rather than solitons, should be considered primary mechanisms in frequency-multiplexed optical computing. With it, an important question related to the system design and performance optimization of frequency-multiplexed optical computing schemes is solved.
Future work should include experimental validation of the regimes of the nonlinear temporal Talbot effect and their transitions. Additionally, studies of more realistic systems that incorporate effects such as higher-order dispersion, loss, and the Raman effect are necessary to better understand and exploit the regimes of the nonlinear temporal Talbot effect in real-world applications. Whereas the current study provides an approximate analysis of the dynamics for low values of the phase modulation of the CW laser, $m\ll 1$, future work should also consider higher values of $m$ and how they fit within the dispersion-related framework of the nonlinear temporal Talbot effect as presented here.
Data availability
@@ Acknowledgments
I would like to say thank you to Michael Böhm (Wismar University of Applied Sciences, Germany), the inventor of Soliton Radiation Beat Analysis, for our long and inspiring conversations about optical solitons.